OpenStudy (anonymous):

two groups entering a museum each paid $70. One group included ten adults and four children. The other included five adults and sixteen children. Find the admission price for children and for adults

OpenStudy (anonymous):

no i need help please

OpenStudy (anonymous):

that is a 16c instead of 6c in the second equation.multiply second equation with 2 then subtract it from the first equation. c will be remaining (the price for children).u can get the value of c.Now put c in any of the equation and find a.

OpenStudy (anonymous):

wait arent you salving for c first and then you and you plug that in for the c in 10a+4(6)=70

OpenStudy (anonymous):

can you show me how you got 6

OpenStudy (shane_b):

Maybe this will make more sense: \[10a+4c=70\]\[5a+16c=70\] Solving the second equation for "a" gives you: \[a=\frac{(70−16c)}{5}\]Now plug this into the other equation and solve for "c": \[10(\frac{(70−16c)}{5})+4c=70\]You should c = 2.5. Now just plug 2.5 back into one of the original equations and solve for "a": \[10(a)+4(2.50)=70\]\[a=6\]

OpenStudy (shane_b):

I swapped around my a's and c's when I typed it out originally...sorry about that.

OpenStudy (shane_b):

Does it make more sense now?